OpenStudy (hulkart):

Which score has a higher relative position, a score of 38 on a test for which = 27 and s = 10, or a score of 262.7 on a test for which = 200 and s = 57? A score of 262.7 Both scores have the same relative position. A score of 38

OpenStudy (perl):

you can use the z score to find relative position

OpenStudy (perl):

$$ \Large { z = \frac{x - \mu}{\sigma } }$$

OpenStudy (perl):

the z score tells you how many standard deviations you are from the mean

OpenStudy (perl):

$$ \Large { z = \frac{x - \mu}{\sigma } \\~\\z = \frac{38 - 27}{10 }=1.1 \\~\\ \\z = \frac{262.7 - 200}{57 }=1.1 }$$